OpenStudy (anonymous):

Siobhan wants to have a beach party in January. Her parents said she can use the garage, which is 36 feet long and 20 feet wide. If she wants to cover the floor with 2 inches of sand, how much sand does she need?

OpenStudy (anonymous):

Volume = length * width * height. Change your units from feet to inches first... so Volume = (36 * 12) * (20 *12) * 2 = 207360 cubic inches if you want this in cubic feet: Volume = 36 * 20 * (1/6) = 120 cubic feet